For a specific model where the consumption function is given as C = 80 + 0.6Y, while Investments are 120, and there is no government purchases and no net exports, a) find the equilibrium level of income. Show your calculations. b) find the level of savings. Show your calculations. c) compare your S from (b) and the given I – what can you say about this economy? d) if, for some reason, output is at the level of 650, what will the level of Unplanned Inventory be? Show your calculations. e) if I drops to 100, what will the effect be on the equilibrium income? – Calculate the new equilibrium level of output. Show your calculations. f) what is the value of the multiplier, α, here – for the original consumption function? Show your calculations.
Answer #1

a) With no government purchases and net exports,

Equilibrium level of income Y = C+I+G+NX (C=consumption,I=investment,G=government purchases,NX=net exports)

or, Y=80+0.6Y+120+0+0

or, 0.4Y=200

or, Y=500

b) Savings S = Y-C = 500-80-(0.6*500)= 500-80-300 = 120

c) Here, S=I=120, that means savings is equal to the investment.

d) Unplanned inventory = 6500-Y =650-500 = 150

e) If I drops to 100,

Y=80+0.6Y+100 (from equation of equilibrium level of income)

or, 0.4Y=180

or, Y=180/0.4 = 450

f) Multiplier = 1/(1-MPC) = 1/(1-0.6) = 1/0.4 = 2.5